# ConfQA: Answer Only If You Are Confident

### v1

@@ 5.1 Fine-tuning setup
We compare our solution with two state-of-the-art solutions. R-Tuning (Zhang et al., 2024) generates its training data by adding a prompt “Are you sure you accurately answered the question based on your internal knowledge?” in the question, and padding “I am sure” or “I am unsure” based on correctness of the generated answer. In the inference, it again appends the prompt and applies post-processing by removing answers with the suffix of “I am unsure”. We used both MMLU (proposed in the paper) and DBPedia for training. IDK (Cheng et al., 2024) requires answer consistency in addition to answer correctness, and we add a consistency requirement of at least four out of five times. We used DBPedia for its fine-tuning for more direct comparison.
Additionally, we considered the following alternatives for ablation study; more alternatives are discussed in Appendix 14.
Implementation and hyper-parameters:
We experimented with Llama-3.1-70B (we observed similar trend on Llama-3.1-8B). We conducted a simple scaling-law study (see Appendix 12) and decided to run one epoch on 3K high quality training data, with a learning rate of 1e-6 and a batch size of 1.
5.2 Results on short-form question answers
Model Corr (Rec) Miss Hall. Fac. Prec F1 Corr (Rec) Miss Hall. Fact. Prec F1
DBpedia (in-domain) IMDB (out-of-domain)
Llama-3.1 52.0 22.0 26.0 26.0 66.7 58.4 44.8 34.2 21.0 23.8 68.1 54.1
Llama-3.1 (D) 47.0 26.8 26.2 20.8 64.2 54.3 40.7 36.2 23.2 17.5 63.7 49.6
R-tuning (MMLU) 24.3 67.3 8.3 16.0 74.5 36.6 28.2 60.5 11.3 16.9 71.3 40.4
R-tuning (DBPedia) 24.5 67.8 7.7 16.8 76.2 37.1 25.3 70.2 4.5 20.8 84.9 39.0
IDK (DBPedia) 17.0 81.5 1.5 15.5 91.9 28.7 22.0 77.0 1.0 21.0 95.7 35.8
ConfQA 49.9 33.5 17.5 31.5 74.0 59.0 43.0 42.0 16.0 27.0 73.0 54.0
ConfQA (D) 31.5 63.3 5.2 26.3 85.9 46.1 32.5 63.3 4.2 28.3 88.6 47.6
SimpleQA (out-of-domain) CRAG (out-of-domain)
Llama3.1 20.0 44.1 35.9 -15.8 35.8 25.7 58.7 15.6 25.7 33.0 69.6 63.7
Llama3.1 (D) 16.8 48.0 35.2 -18.4 32.3 22.1 57.5 22.3 20.2 37.2 73.9 64.7
R-tuning (MMLU) 20.3 38.0 41.7 -21.4 32.8 25.1 57.8 17.1 25.1 32.7 69.7 63.2
R-tuning (DBPedia) 3.7 83.3 13.0 -9.3 22.0 6.3 31.6 55.0 13.4 18.2 70.2 43.6
IDK (DBPedia) 0.6 99.1 0.2 0.4 73.0 1.2 20.7 78.2 1.1 19.6 95.0 34.0
ConfQA 17.3 55.8 26.8 -9.5 39.2 24.0 57.0 19.6 23.4 33.6 70.9 63.0
ConfQA (D) 4.9 93.1 2.1 2.8 70.2 9.1 39.4 56.2 4.4 35.0 90.0 54.8
Table 2: Overall factuality improvement on short-form benchmarks; ConfQA can reduce hallucination to below 5% with the dampener prompt. All numbers are in percentage (%).
Overall results:
@@ 5.2 Results on short-form question answers
Table 2 presents our overall results, and shows interesting effect of the dampener prompt. First, without fine-tuning, the effect of the dampener is inconsistent. For all benchmarks, the dampener increased the percentage of missing answers by 2-7%. However, it (correctly) dampens hallucinations on CRAG, but dampens correct answers and thus reduced the factuality on Head-to-tail and SimpleQA. This is not surprising since the LLM confidence is not well calibrated.
@@
Second, after the fine-tuning, ConfQA improves precision by up to 40% and improves factuality by up to 20%. Without the dampener, we observe mild reduction of the hallucinations, accompanied with minor sacrifice on correct answers, leading to factuality improvement up to 6%. With the dampener, the LLM is much more conservative and the hallucination rate drops to below 5% on all benchmarks. As a side effect, correctness also drops significantly; for example, since SimpleQA focuses on nuanced facts, after the finetuning we observe nearly zero correctness. Still, the factuality increases by up to 18% compared to the baseline.
@@
Third, R-tuning mostly has higher precision and lower hallucination, especially if trained on DBPedia. However, we also observe much lower correctness. We suspect this is because when the model gives a wrong answer, it feeds ground truths as additional knowledge and causes confusion; as we will discuss soon, we observe similar trend for Fact-feeding, with a similar flavor. We also observe better results when trained on DBPedia than on MMLU, which mixes factual statements and skills and can cause confusion.
@@
Fifth, even though the training data are generated only from DBPedia, ConfQA behavior changes on the other datasets as well, showing amazing generalization.
Finally, we note that although the factuality metric increases most of the time, the F1-metric stays flat or drops, as the latter is more lenient to hallucinations as discussed in Appendix 10.
@@
Corr Miss Hall Fac Corr Miss Hall Fac Corr Miss Hall Fac Corr Miss Hall Fac
Llama-3.1 47.0 26.8 26.2 20.8 40.7 36.2 23.2 17.5 16.8 48.0 35.2 -18.4 57.5 22.3 20.2 37.2
@@
* Fact-feeding 20.7 76.7 2.7 18.0 25.5 70.7 3.8 21.7 2.5 94.7 2.8 -0.3 22.4 74.5 3.1 19.3
Table 3: Ablation study when applying dampening in inference, showing effectiveness of our fine-tuned model. All numbers are in percentage (%) and full results in Table 8.
@@
We now compare ConfQA with the several alternatives, as shown in Table 3 (full comparison in Table 8 in Appendix 14). MMLU-as-source obtains low hallucination (below 2%) but also significantly lower correctness. We suspect this is because MMLU contains a diverse set of tasks, reducing overall confidence of the model.
On the other hand, No-dampener increases correctness and reduces missing rate, but can increase hallucinations, showing the important role of the dampener in training as well. GT-as-label achieves the highest correctness and lowest missing rate, but becomes over-confident to hallucinate (hallucination rate can reach 70%). This is consistent with observations in previous work (Lin et al., 2024; Gekhman et al., 2024) that feeding facts in the SFT-stage can teach LLMs to hallucinate.
Interestingly, Fact-feeding combines our ConfQA fine-tuning data with Tulu facts, drops hallucinations but also correctness, similar to R-tuning. We suspect this is because what our training data teach the LLM (saying unsure) is of different purpose from what the extra Tulu facts teach the LLM (feeding knowledge), when mixed together can offset each other and cause confusion.
@@
Figure 3: ConfQA suppresses more on long-tail facts than popular facts.
@@
Finally, we show in Figure 3 the distribution of correct, missing, and incorrect answers for entities of different popularity, before and after fine-tuning, with and without dampening. It confirms that fine-tuning suppresses hallucinations, and the dampener prompt further reduces hallucinations. Additionally, it shows ConfQA suppresses more on long-tail facts, where it lacks confidence.
@@ 5.3 Results on other benchmarks
Figure 4: RAG invocation architecture.
We also evaluate ConfQA on the standard MMLU benchmark. Table 5 shows that the scores are mostly similar to the baseline.
Summary:
Our experiments validated the effectiveness of our ConfQA fine-tuning strategy. If one aims to eliminate hallucinations, we recommend applying ConfQA with the dampener; if one aims to maintain the correctness and meanwhile reduce hallucinations as much as possible, we recommend ConfQA without the dampener.
6 Q3. What Is the Optimal Strategy for Triggering RAG?
Since our fine-tuned model can reduce hallucinations to nearly zero, we can invoke the RAG pipeline when it says unsure, and rely on the LLM-generated answer otherwise. Figure 4 depicts a RAG invocation architecture. The system always answers a dynamic question that inquires changing information through the RAG pipeline. When a static question arrives, the system invokes LLM generation and the RAG pipeline in parallel. When LLM generates a non-missing answer, the system early-stops the RAG pipeline and outputs the answer; otherwise, the system waits and outputs the answer from the RAG pipeline.
Calculations based on Table 2 shows that our framework enables potential accuracy gains to beyond 95%, while reducing unnecessary external retrievals by over 30%. We next evaluate QA quality and latency through a real RAG implementation, which invokes search APIs (Bing API and Knowledge Graph API) for retrieval, and passes the retrieved content to Llama-3.1-70B to generate the responses. Table 6 reports the end-to-end QA accuracy and latency for our proposed RAG architecture, and compares it with not invoking RAG and invoking RAG everywhere. ConfQA-triggered RAG obtains similar quality to RAG-everywhere, but cut latency by 300ms P50 and 400ms P90 for CRAG. The latency-cut is less pronounced on SimpleQA since RAG should be triggered for the majority of the questions. The quality improvement is small on CRAG because it contains a lot of complex questions requiring reasoning over retrieval results, which our simple RAG implementation does not excel.
Model SimpleQA CRAG
Upper Corr Miss Hall Fac L-P50 L-P90 Upper Corr Miss Hall Fac L-P50 L-P90
Table 6: ConfQA-based RAG invocation achieves similar quality to RAG-everywhere, while cutting latency.
@@ 7 Limitations
Our experiments focus on SFT, and we leave DPO-based fine-tuning for future work. We conducted training on the Llama-3.1 model of different sizes; it would be interesting to experiment with other open sourced models and later versions of models. We empirically compared DBPedia and MMLU, where the former contains only simple factual questions, and the latter contains questions ranging from factual to reasoning. A comprehensive study regarding the effect of sources with different coverage in this spectrum would improve the understanding. We can extend our learnings beyond factual statements, for math, coding, reasoning etc. Lastly, SFT requires the access to LLM itself for fine tuning, and thus restrict the application of the proposed framework to proprietary LLMs only accessible through APIs.
@@ 8 Conclusion
Recent studies have shown that LLMs acquire substantial knowledge during pre-training, and that introducing new knowledge during post-training can often increase hallucinations (Lin et al., 2024; Gekhman et al., 2024). Rather than injecting additional knowledge, the Dual Neural Knowledge (DualKnowl) framework we propose trains LLMs to identify and withhold low-confidence factual claims, deferring to RAG in such cases to improve factuality. Through a comprehensive set of experiments, we show that 1) LLMs tend to be over-confident on what they know; 2) we can fine-tune LLM to refrain from generating inconfident factual statements and thus reducing the hallucination rate to below 5%; and 3) using this fine-tuned model for RAG-triggering can reach similar answer accuracy as RAG-everywhere, whereas reducing RAG retrievals by over 30% to save cost. Our proposed framework naturally integrates internal neural knowledge with external symbolic knowledge, allowing for improvements in both factuality and latency as LLMs develop richer internal knowledge and more effective RAG capabilities.

### v2

@@ 5.1 Alternative Solutions
We compare ConfQA with two baseline solutions: LLM without dampener and LLM with dampener in the inference. In addition, we implemented two state-of-the-art solutions. R-Tuning (Zhang et al., 2024a) generates its training data by adding a prompt “Are you sure you accurately answered the question based on your internal knowledge?” in the question, and padding “I am sure” or “I am unsure” based on correctness of the generated answer. In the inference, it again appends the prompt and applies post-processing by removing answers with the suffix of “I am unsure”. We used both MMLU (Hendrycks et al., 2021) (proposed in the paper) and DBPedia for training. IDK (Cheng et al., 2024) requires answer consistency in addition to answer correctness, and we add a consistency requirement of at least four out of five times. We used DBPedia for its fine-tuning for more direct comparison.
5.2 Effectiveness of ConfQA (RQ2)
Model Corr Miss Incor Fac. Ceiling Corr Ceiling Fac. Tri. Prec Tri. Rec Fmsr
DBpedia (in-domain)
Llama-3.1 52.0 22.0 26.0 26.0 74.0 48.0 - - -
Llama-3.1 Dampen 47.0 26.8 26.2 20.8 73.8 47.6 100.0 50.6 67.2
R-tuning (MMLU) 24.3 67.3 8.3 16.0 91.6 83.3 71.3 85.3 77.7
R-tuning (DBPedia) 24.5 67.8 7.7 16.8 92.3 84.6 70.8 86.2 77.7
IDK (DBPedia) 17.0 81.5 1.5 15.5 98.5 97.0 58.9 97.0 73.3
ConfQA 31.5 63.3 5.2 26.3 94.8 89.6 75.8 90.2 82.4
IMDB (out-of-domain)
Llama-3.1 44.8 34.2 21.0 23.8 79.0 58.0 - - -
Llama-3.1 Dampen 40.7 36.2 23.2 17.5 76.9 53.7 100.0 60.9 75.7
R-tuning (MMLU) 28.2 60.5 11.3 16.9 88.7 77.4 91.2 83.0 86.9
R-tuning (DBPedia) 25.3 70.2 4.5 20.8 95.5 91.0 78.6 92.5 85.0
IDK (DBPedia) 22.0 77.0 1.0 21.0 99.0 98.0 71.7 98.2 82.9
ConfQA 32.5 63.3 4.2 28.3 95.8 91.6 87.2 92.9 90.0
SimpleQA (out-of-domain)
Llama3.1 20.0 44.1 35.9 -15.8 64.1 28.2 - - -
Llama3.1 Dampen 16.8 48.0 35.2 -18.4 64.8 29.6 100.0 57.7 73.2
R-tuning (MMLU) 20.3 38.0 41.7 -21.4 58.3 16.6 100.0 47.7 64.6
R-tuning (DBPedia) 3.7 83.3 13.0 -9.3 87.0 74.0 96.0 86.0 90.8
IDK (DBPedia) 0.6 99.1 0.2 0.4 99.7 99.5 80.7 99.8 89.2
ConfQA 4.9 93.1 2.1 2.8 98.0 95.9 85.9 97.4 91.3
CRAG (out-of-domain)
Llama3.1 58.7 15.6 25.7 33.0 74.3 48.6 - - -
Llama3.1 Dampen 57.5 22.3 20.2 37.2 79.8 59.6 100.0 52.5 68.8
R-tuning (MMLU) 57.8 17.1 25.1 32.7 74.9 49.8 100.0 40.5 57.7
R-tuning (DBPedia) 31.6 55.0 13.4 18.2 86.6 60.9 75.1 75.5 75.3
IDK (DBPedia) 20.7 78.2 1.1 19.6 98.9 97.8 52.8 97.4 68.5
ConfQA 39.4 56.2 4.4 35.0 95.6 91.2 73.5 90.4 81.1
Table 2: Overall factuality and triggering Fmsr improvement on short-form benchmarks; ConfQA can reduce hallucination to below 5%. ConfRAG achieves the best Fmsr over all methods among all benchmarks. The optimal Fmsr, Tri. Prec, Tri. Rec, Factuality and Incorrect rate are shown in bold. The second best Incorrect rate are shown in italic. Dash - indicates the metrics are not valid as as we use Llama-3.1 triggering as ground truth. All numbers are in percentage (%).
@@ 5.2 Effectiveness of ConfQA (RQ2)
Table 2 left part presents the answer quality from ConfQA. First, baseline methods are hallucination prone. Without fine-tuning, the effect of the dampener is inconsistent. For all benchmarks, the dampener increased the percentage of missing answers by 2-7%. However, it (correctly) dampens hallucinations on CRAG, but dampens correct answers and thus reduced the factuality on Head-to-tail (DBPedia) and SimpleQA. This is not surprising since the LLM confidence is not well calibrated.
@@
Second, the fine-tuned ConfQA improves factuality by up to 20%, and the hallucination rate drops to below 5% on all benchmarks. As a side effect, correctness also drops; for example, since SimpleQA focuses on nuanced facts, after the finetuning we observe nearly zero correctness. Still, there is much more dropping on hallucniated answers than on correct answers, and thus the factuality increases across benchmarks.
@@
Third, R-tuning mostly has lower hallucination, especially if trained on DBPedia. However, we also observe much lower correctness. We suspect this is because when the model gives a wrong answer, the training data feed ground truths as additional knowledge and causes confusion. We also observe stronger performance when trained on DBPedia than on MMLU, as MMLU mixes facts with reasoning skills and can introduce ambiguity, supporting our hypothesis that atomic facts yield better training examples.
@@
Finally, even though the training data are generated only from DBPedia, ConfQA behavior changes on the other datasets as well, showing amazing generalization.
Model SimpleQA CRAG
Upper Corr Miss Incor Fac L-P50 L-P90 Upper Corr Miss Incor Fac L-P50 L-P90
Table 3: ConfQA-based RAG invocation achieves similar quality to RAG-everywhere, while cutting latency. Upper: upper bound of percentage of correct. L-P50: P50 latency and L-P90: P90 latency.
5.3 Effectiveness of ConfRAG (RQ3)
Table 2 also compares the effectiveness of triggering using different QA solutions, showing that ConfRAG has the highest triggering F-measure for each datasets, enables potential truthfulness gains to beyond 95%, while reducing unnecessary external retrievals by 5-19%.
We next evaluate QA quality and latency through a real RAG implementation, which invokes search APIs (Bing API and Knowledge Graph API) for retrieval, and passes the retrieved content to Llama-3.1-70B to generate the responses. Table 3 reports the end-to-end QA accuracy and latency for our proposed RAG architecture, and compares it with not invoking RAG and invoking RAG everywhere. ConfQA-triggered RAG obtains similar quality to triggering RAG everywhere, but reduced latency by 600ms P50 and 800ms P90 for CRAG. The SimpleQA benchmark requires triggering RAG for the majority of the questions to achieve high quality; the latency improvement is less pronounced but our triggering does not regress quality compared to RAG-everywhere. The CRAG benchmark contains a lot of complex questions requiring reasoning over retrieval results; our simple RAG implementation does not excel, but still improves on factuality.
@@ 5.3 Effectiveness of ConfRAG (RQ3)
Corr Miss Incor Fac Corr Miss Incor Fac Corr Miss Incor Fac Corr Miss Incor Fac
Llama-3.1 Dampen 47.0 26.8 26.2 20.8 40.7 36.2 23.2 17.5 16.8 48.0 35.2 -18.4 57.5 22.3 20.2 37.2
@@
Table 4: Ablation study for ConfQA when applying dampener in inference for all models (unless explicitly says no), showing effectiveness of our fine-tuned model. Optimal Correct, Factuality and two best Incorrect are shown in bold. All numbers are in percentage (%) and full results in Table 8.
5.4 Deep dive on ConfQA
@@
We now compare ConfQA with the several alternatives, as shown in Table 4 (full comparison in Table 8 in Appendix 16). First, without the dampener in inference, we observe minor sacrifice or even slight increase on correct answers, but also just mild reduction of the hallucinations. On the other hand, without the dampener in training, we observe increased correctness and reduced missing rate, but also increased hallucinations than ConfQA and thus lower factuality, showing the important role of the dampener in training as well.
Use MMLU, instead of DBPedia, to generate training data obtains low hallucination (below 2%) but also significantly lower correctness. We suspect this is because MMLU contains a diverse set of tasks, reducing overall confidence of the model. GT-as-label achieves high correctness and lowest missing rate, but becomes over-confident to hallucinate (hallucination rate can reach 70%). This is consistent with observations in previous work (Lin et al., 2024; Gekhman et al., 2024) that feeding facts in the SFT-stage can teach LLMs to hallucinate.
@@
Figure 4: ConfQA suppresses more (gives more missing answers) on less popular entities (missing rate for Tail is larger than Torso, whose missing rate is larger than Head).
@@
Finally, we show in Figure 4 the distribution of correct, missing, and incorrect answers for entities of different popularity (Head, Torso, Tail), before and after fine-tuning, with and without dampening. It confirms that fine-tuning suppresses hallucinations, and the dampener prompt further reduces hallucinations. Additionally, it shows ConfQA suppresses more on long-tail facts, where it lacks confidence.
6 Discussions
In addition to triggering, we also examined ConfQA on other benchmarks, to understand whether it can apply to suppress hallucination when access to retrieval corpus is unavailable, such as in most speech-in speech-out systems (Xie and Wu, 2024; Nguyen et al., 2022; Zhang et al., 2024b). We wish to reduce hallucinations on factual statements, without regressing performance on general knowledge and problem-solving tasks.
@@ 6 Discussions
We thus consider benchmarks in two other categories: 1) Long-form factuality benchmarks, where answers are expected to be long and contain multiple factual statements. We use the automatic evaluation metric, VeriScore (Song et al., 2024), which computes precision, recall, and F1-score. We set the minimum number of facts required for perfect recall based on the median number of extracted claims per dataset, using their fine-tuned models for claim extraction and verification. 2) General knowledge benchmarks, which focuses on general knowledge and reasoning in diverse disciplines. MMLU provides ground truths for the multi-choice questions. The score is computed as the percentage of correctly answered questions, as a weighted average among the 57 diverse subjects. Again, Table 1 summarizes the benchmarks and Appendix 10 gives details.
We also evaluate ConfQA on the standard MMLU benchmark, and also do not apply the dampener prompt. Table 6 shows that the scores are mostly similar to the baseline.
Together with the results presented in Section 5, the experiments suggest another potential application of ConfQA when RAG is not an option. If one aims to maintain the correctness and meanwhile reduce hallucinations as much as possible, we recommend ConfQA without the dampener; if one emphasizes hallucination elimination, we recommend applying ConfQA with the dampener.
@@ 7 Limitations
Our experiments focus on SFT, and we leave DPO-based fine-tuning for future work. We empirically compared DBPedia and MMLU, where the former contains only simple factual questions, and the latter contains questions ranging from factual to reasoning. A comprehensive study regarding the effect of sources with different coverage in this spectrum would improve the understanding. We can extend our learning beyond factual statements, for math, coding, reasoning etc. Lastly, SFT requires the access to LLM itself for fine tuning, and thus restrict the application of the proposed framework to proprietary LLMs only accessible through APIs.
@@ 8 Conclusion
Recent studies have shown that LLMs acquire substantial knowledge during pre-training, and that introducing new knowledge during post-training can often increase hallucinations (Lin et al., 2024; Gekhman et al., 2024). Rather than injecting additional knowledge, the ConfRAG framework we propose trains LLMs to identify and withhold low-confidence factual claims, deferring to RAG in such cases to improve factuality. Through a comprehensive set of experiments, we show that 1) LLMs tend to be over-confident on what they know; 2) we can fine-tune LLM to refrain from generating inconfident factual statements and thus reducing the hallucination rate to below 5%; and 3) using this fine-tuned model for RAG-triggering can reach similar answer accuracy as RAG-everywhere, whereas reducing RAG retrievals by over 5-19% to save cost. Our proposed framework naturally integrates internal neural knowledge with external symbolic knowledge, allowing for improvements in both factuality and latency as LLMs develop richer internal knowledge and more effective RAG capabilities.